OpenStudy (anonymous):

Help please ! The turnout at a concert depends on the number of posters put up around town, so the turnout is a function of the number of posters that are hung. If this function is called T(p) use the formula below to compute T(9), the turnout at a concert that was advertised using 9 posters. T(p)=60p-p^2

OpenStudy (anonymous):

\[T(9)=60\cdot 9-9^2\]

OpenStudy (nurali):

T(p) = 60p - p2 for T(9), replace p with 9 T(9) = 60*9 - 9*9 T(9) = 540 - 81